Stacy loaned Juan $1040
at an interest rate of 15%
for 226
days. How much will Juan pay Stacy at the end of 226
days? Round your answer to the nearest cent. Note: Assume 360
days in a year and 30
days in a month.

Answer by mananth(16946)

Stacy loaned Juan $1040
at an interest rate of 15%
for 226 days. How much will Juan pay Stacy at the end of 226
days? Round your answer to the nearest cent. Note: Assume 360
days in a year and 30
days in a month.
SI = P*n*R
SI is simple interest
P is the principal
n is the number of years
r=15%=0.15
SI= 1040*0.628*0.15 =97.97

1040+97.97=1137.97

Therefore, at the end of 226 days, Juan will pay Stacy approximately $1137.97